OpenStudy (anonymous):

write the equation a slope-intercept form of the line that slope of 2 and contains the point (1,1) help

OpenStudy (amistre64):

y = mx-mPx+Py is my remembered format

OpenStudy (cwrw238):

use the general form y-y1 = m(x-x1) where m = slope and (x1,y1) is a point on the line

OpenStudy (amistre64):

m=2, (Px,Py) -> (1,1)